OpenStudy (anonymous):

cos2a = 1 - 2sin^2 a TRUE or FALSE

OpenStudy (anonymous):

True...It's also an identity.

OpenStudy (anonymous):

Just making sure, it is true! :) Thanks

OpenStudy (anonymous):

\[\bf \cos(2a)=1-2\sin^2(a) \ \leftarrow TRUE\]

OpenStudy (anonymous):

COme on bro... I know you knew that! @ilfy214

OpenStudy (anonymous):

Lolz! I must make sure its correct!

OpenStudy (anonymous):

:-)